# The first catalogue of candidate white dwarf-main sequence binaries in open star clusters: A new window into common envelope evolution

### v1

@@ 6.3.2 Main-Sequence+Main-Sequence Binaries
We find that the MS+MS binaries almost exactly match the Pickles curves shown in Figure 5 for both extinction values. For binaries with hotter MS stars (e.g. F-type) and cooler companions (e.g. M-type), the relative flux ratios are such that the hotter star dominates the photometry, appearing as a single MS star. On the other hand, while similar mass combinations appear brighter in the Gaia CMD, their colours are only minimally impacted. When run through our SVM model, zero MS+MS model binaries are classified as high-probability WD+MS systems for both the $A_{V}=0.0$ and $A_{V}=3.0$ cases, with the highest probability system being classified at $P=0.02$ (since these Pickles models were included in our training sample, it is unsurprising that they almost perfectly align with the MS photometry). Hence, we expect very little to no contamination from MS+MS binaries in our our catalogue.
@@ 6.3.3 Rapidly Rotating M-dwarfs
A significant fraction of low-mass stars are highly (or even fully) convective. Strong convection can generate large magnetic fields in stellar chromospheres, causing immense stellar activity in the form of eruptions or flares (Vallée, 2003). The General Catalogue of Variable Stars (Samus’ et al., 2017) lists more than twenty sub-classes of eruptive variables, broadly categorized by their spectral type, flare frequency/regularity and astrophysical origin. Of particular interest are flaring (active) M-dwarf stars, which predominantly erupt in the near-ultraviolet portion of the electromagnetic spectrum and can thus manifest as blue colour excesses in M-dwarf photometry if a flare was active when observations were obtained (Welsh et al., 2007; Li et al., 2024). While the exact mechanisms that drive stellar flaring are still an open question, observational studies of chromospheric activity in M-dwarfs are extensive (e.g. Davenport, 2016; Kado-Fong et al., 2016; Newton et al., 2017; Yang et al., 2017; Medina et al., 2020; Feinstein et al., 2020; Rodríguez Martínez et al., 2020; Rekhi et al., 2023).
@@
Since flares in M-dwarfs can contribute excess blue flux to otherwise red optical photometry, we consider how active M-dwarfs are classified with our SVM. Newton et al. (2017) provide a catalogue of M-dwarfs with detailed activity indicators, however these stars are all located within the Solar Neighbourhood and are thus too bright to have reliable (unsaturated) photometric measurements in Pan-STARRS1. Thus, we consider the Kado-Fong et al. (2016) catalogue of M-dwarfs observed in the Pan-STARRS1 Medium-Deep Survey, since these M-dwarfs are located at further distances. Kado-Fong et al. (2016) identify 270 rotating M-dwarfs, measuring periods of 0.7 days $<P<$ 130 days. While no activity indicators are explicitly included in this catalogue, Newton et al. (2017) observe that almost all M-dwarfs with rotational periods of $P<10$ days are active. After cross-matching this catalogue with 2MASS, Gaia DR3 and Pan-STARRS1, we find 126 M-dwarfs in the Kado-Fong et al. (2016) sample with $P<10$ days. 110 of these stars adhere to the same photometric and kinematic constraints described in Section 2.2.2 and 2.2.3. We adopt this sample of rapidly rotating M-dwarfs as a proxy for the expected properties of active M-dwarfs.
@@
From Figure 12, we see that active M-dwarfs (red triangles) lie directly along the MS of our training data in both a CMD and all colour-colour spaces. Thus, we find that only three systems ($3\%$) are classified as high-probabilty WD+MS candidates with our SVM. However, despite these low numbers, the WD+MS candidates that lie close the ZAMS (approximately 25-30% of our final sample) do share some photometric properties with active M-dwarf stars. For these systems, we conclude that some contamination is possible (especially since single M-dwarfs are expected to be more prevalent than WD+MS binaries) and spectroscopy is required to distinguish binarity from potential flaring. In particular, we note (i) Pleiades-c1 is located in this region of parameter space, but is a spectroscopically confirmed binary system and (ii) both Pleiades-c2 and Pleiades-c3 were previously classified as flaring M-dwarfs (Section 6.2.2). However, also note that the presence of M-dwarf flares is not mutually exclusive with a potential WD+MS binary, as post-CE systems with active M-dwarf secondary stars have been identified (e.g. Muirhead et al., 2013).
@@ 6.3.4 Single WDs
Depending on system age, single WDs can have effective temperatures between $10,000K<T_{eff}<40,000K$ (Eisenstein et al., 2006; Gentile Fusillo et al., 2021). Hence, it is possible that older, cooler and redder WDs could share similar photometry to a WD+MS binary. To investigate this potential overlap, we consider the Gentile Fusillo et al. (2021) catalogue of WDs identified in Gaia Early Data Release 3. While the full sample contains $\sim 359,000$ WDs, we find that upon cross-matching, only 88 of these systems have 2MASS data.
In Figure 12, we observe that these 88 single WDs fall bluewards of the previously investigated populations in both a Gaia CMD and all colour-colour spaces. While single WDs appear to occupy extensions of the WD+MS binary population in each colour-colour space, overlap with WD+MS binaries is present especially when 2MASS photometry is included. Due to this overlap, $38\%$ of the WDs in this catalogue are classified by our SVM as probable WD+MS binaries. However, none of our identified WD+MS candidates visually overlap with the single WDs in any parameter space, making it unlikely that any of our candidates are actually single WDs.
Figure 13: The proposed effect of a CE natal velocity kick on the spatial position and kinematics of candidate NGC1342-c1. Left: Spatial location of NGC1342-c1 (star) relative to high-probability cluster stars from Cantat-Gaudin et al. (2020) (gray). Each circle represents the distance NGC1342-c1 would be located if it received a natal kick for various kick magnitudes and travel times. Right: Similar to the left panel, however with proper motion. Each circle here represents the change in NGC1342-c1’s proper motion ($\Delta\mu$) upon receiving a velocity kick completely imaprted into the tangential direction (for simplicity).
@@ 6.4 Implications for Natal Kicks
Upon examining the kinematics of the candidate WD+MS systems in our catalogue, we find that while their overall proper motions are consistent with their host clusters, the spatial locations for more than half of the candidates are largely offset (see Section 4.2). Thus, we explore whether these spatial offsets could be explained by invoking a natal kick. As a fiducial example, we consider the WD+MS binary candidate NGC1342-c1. NGC1342-c1 has a proper motion consistent with being a member of NGC 1342 ($\chi^{2}=1.03$) but a spatial location inconsistent with the cluster ($\chi^{2}=112.41$) as computed in Section 4.2.
@@
While specific cluster proper motions, distances and escape velocities determine the magnitude to which a natal kick might effect a post-CE system, this analysis shows that a post-CE system can be spatially offset from a cluster despite having a consistent proper motion. Hence, we conclude that it is possible for post-CE systems to migrate far from their birth cluster if a velocity kick is received. Importantly, we emphasize that $78\%$ of the OCs searched in this study revealed no WD+MS candidates. It is almost certain that there exist more WD+MS post-CE cluster binaries than presented in this catalogue, however they are likely located outside the search radii chosen in this study. While extending search radii might help identify more WD+MS candidates, finding binaries that have travelled far beyond their host cluster is challenging. Although chemical tagging and dynamical simulations are good tools for associating escaped stars and binaries with individual star clusters (e.g. Grondin et al., 2023, 2024), ample chemistry and radial velocities would be required for definitive associations of these escaped systems.
@@ 7 Summary and Conclusions
In this paper, we search for WD+MS binaries in 299 OCs in the Cantat-Gaudin et al. (2020) catalogue. Using photometry from Gaia DR3, 2MASS and Pan-STARRS1, we train a support vector machine model to identify systems that are photometrically similar to a sample of SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010). Throughout, we design both our training sample and method to try to minimize contamination (and thus sacrifice some amount of completeness) in order to identify a set of high-probability candidates, which can then be followed-up for confirmation and characterization. While the focus of this study is identifying WD+MS binaries in clusters, we emphasize that this methodology could be utilized for photometric identification of a variety of other astrophysical sources. Figure 1 highlights the main steps in our identification method. Our main results are summarized below.
@@
    This catalogue contains 52 candidate WD+MS binaries in 38 OCs which were identified with high-probability ($P>0.9$) from our SVM classification model and subsequently vetted for data quality (Figure 8). Generally, our catalogue contains systems that are redder than the overall sample of WD+MS binaries presented in Rebassa-Mansergas et al. (2010). This is in line with other WD+MS systems identified in Gaia (Inight et al., 2021; Rebassa-Mansergas et al., 2021). 78% of the 299 OCs searched had zero high-probability WD+MS binaries identified by our method.
@@
    We present follow-up spectroscopy from the Lick and Gemini Observatories for three example candidates that span a range of Gaia $BP-RP$ colours (Figure 10). From this spectroscopy alone, it is clear that a variety of systems are present in this catalogue. For instance, two objects (which represent both the bluest and reddest systems in our sample) are well-matched to M-type MS stellar models at red wavelengths, but show evidence of broad Balmer lines and excess flux at blue wavelengths—as would be expected for a WD companion. In contrast, the third object shows no clear signatures of a WD in its optical spectrum. It may represent either a case with a limited flux contribution from the WD in the optical regime, or contamination in our sample from an active (flaring) M-dwarf. SEDs for all WD+MS candidate binaries are similar to the three examples shown. Hence, we expect that this catalogue spans a wide range of WD and MS configurations.
@@
    Since we solely use photometry to identify WD+MS candidates, we investigate the possibility that other astrophysical sources are present in our catalogue. Specifically, we run populations of cataclysmic variables (Inight et al., 2023), MS+MS binaries (Pickles, 1998), rapidly rotating (active) M-dwarfs (Kado-Fong et al., 2016) and single WDs (Gentile Fusillo et al., 2021) through our SVM. Based on this and the locations of our candidate systems in the Gaia CMD and various colour-colour spaces, we conclude that we are unlikely to have any significant contamination from single WDs or MS+MS binaries. In contrast, our sample may contain CVs in addition to detached WD+MS binaries, and we may have some contamination from young, active M-dwarfs in the $\sim 25-30\%$ of our sample that lies close the the ZAMS. Ultimately, follow-up spectroscopy will be required to fully confirm the nature of each candidate.
@@
Ultimately, this catalogue is a necessary first step in a larger effort to provide observational constraints on the CE phase and detailed characterization of a subset of candidates (e.g. Alessi12-c1, Pleiades-c1, etc.) identified in this study is currently underway. Once a larger sample of post-CE binaries in clusters is confirmed, pre-CE progenitor masses can be estimated, leading to a one-to-one mapping between the initial and final masses of systems that have undergone a CE event (similar to how the IFMR for isolated WDs is determined). With these observational benchmarks, this sample will aid in efforts to unlock important new insights into one of the most uncertain phases of binary evolution.
@@ Acknowledgements
The authors thank Max Moe for providing insightful discussion about potential contaminant sources that aided in an improved understanding of the WD+MS candidates in our catalogue. The authors also thank Alexander Laroche for useful input in validating our machine learning model and Fraser Evans and Phil Van-Lane for valuable edits in the manuscript. Helpful conversations with Samantha Berek, Victor Chan, Mairead Heiger, Marten van Kerkwijk, Gustavo Medina-Toledo and Jeremy Webb also greatly improved this research.
@@ 8 Data Availability
A machine readable version of this catalogue will be uploaded to Vizier upon acceptance of this manuscript. Until then, data will be shared upon reasonable request.
@@ Appendix A Photometric Search Examples
Here, we present results from the SVM classification for two example OCs: Alessi 12 and NGC 2682. As seen in both Figures 14 and 15, the majority of sources in the Alessi 12 and NGC 2682 stellar samples lie along/near the ZAMS in both sample colour-colour spaces and a Gaia CMD. While the SVM does not identify any probable WD+MS candidate binaries in NGC 2682, one high-probability source ($P>0.9$) and two medium-probability sources ($P>0.5$) are found in Alessi 12.
@@ Appendix B Photometric Errors of High-Probability Candidates
Here, we present the photometric errors of all 52 high-probability candidate WD+MS binaries in this catalogue. Overall, the photometry (including error bars) spans the full WD+MS sample, with only small overlap with the ZAMS. We note that the largest photmetric errors arise from 2MASS $J$ and $K$ bands.
Figure 16: Photometric errors of the high-probability candidate WD+MS binaries in this catalogue.
@@ Appendix C Spectral Energy Distributions
As described in Section 3.7, we perform additional vetting of our high-probability candidates through a visual inspection of candidate SEDs. Below, we present SEDs for all 52 high-probability candidate WD+MS systems in our catalogue. We include both Pan-STARRS $g$, $r$, $i$, $z$ and $y$ band along with 2MASS $J$, $H$ and $K$ band photometry. To guide the eye, we have included five MS Pickles models (Pickles, 1998) spanning K0V to M5V spectral types. However, we emphasize that the observational SEDs have not been corrected for Galactic extinction.
Figure 17: Spectral energy distributions of the high-probability candidate WD+MS binaries in this catalogue.

### v2

@@ 6.3.2 Main-Sequence+Main-Sequence Binaries
We find that the MS+MS binaries almost exactly match the Pickles curves shown in Figure 5 for both extinction values. For binaries with hotter MS stars (e.g. F-type) and cooler companions (e.g. M-type), the relative flux ratios are such that the hotter star dominates the photometry, appearing as a single MS star. On the other hand, while similar mass combinations appear brighter in the Gaia CMD, their colors are only minimally impacted. When run through our SVM model, zero MS+MS model binaries are classified as high-probability WD+MS systems for both the $A_{V}=0.0$ and $A_{V}=3.0$ cases, with the highest probability system being classified at $P=0.02$ (since these Pickles models were included in our training sample, it is unsurprising that they almost perfectly align with the MS photometry). Hence, we expect very little to no contamination from MS+MS binaries in our catalog.
@@ 6.3.3 Rapidly Rotating M-dwarfs
A significant fraction of low-mass stars are highly (or even fully) convective. Strong convection can generate large magnetic fields in stellar chromospheres, causing immense stellar activity in the form of eruptions or flares (Vallée, 2003). The General catalog of Variable Stars (Samus’ et al., 2017) lists more than twenty sub-classes of eruptive variables, broadly categorized by their spectral type, flare frequency/regularity and astrophysical origin. Of particular interest are flaring (active) M-dwarf stars, which predominantly erupt in the near-ultraviolet portion of the electromagnetic spectrum and can thus manifest as blue color excesses in M-dwarf photometry if a flare was active when observations were obtained (Welsh et al., 2007; Li et al., 2024). While the exact mechanisms that drive stellar flaring are still an open question, observational studies of chromospheric activity in M-dwarfs are extensive (e.g. Davenport, 2016; Kado-Fong et al., 2016; Newton et al., 2017; Yang et al., 2017; Medina et al., 2020; Feinstein et al., 2020; Rodríguez Martínez et al., 2020; Rekhi et al., 2023).
@@
Since flares in M-dwarfs can contribute excess blue flux to otherwise red optical photometry, we consider how active M-dwarfs are classified with our SVM. Newton et al. (2017) provide a catalog of M-dwarfs with detailed activity indicators, however these stars are all located within the Solar Neighbourhood and are thus too bright to have reliable (unsaturated) photometric measurements in Pan-STARRS1. Thus, we consider the Kado-Fong et al. (2016) catalog of M-dwarfs observed in the Pan-STARRS1 Medium-Deep Survey, since these M-dwarfs are located at further distances. Kado-Fong et al. (2016) identify 270 rotating M-dwarfs, measuring periods of 0.7 days $<P<$ 130 days. While no activity indicators are explicitly included in this catalog, Newton et al. (2017) observe that almost all M-dwarfs with rotational periods of $P<10$ days are active. After cross-matching this catalog with 2MASS, Gaia DR3 and Pan-STARRS1, we find 126 M-dwarfs in the Kado-Fong et al. (2016) sample with $P<10$ days. 110 of these stars adhere to the same photometric and kinematic constraints described in Section 2.2.2 and 2.2.3. We adopt this sample of rapidly rotating M-dwarfs as a proxy for the expected properties of active M-dwarfs.
@@
From Figure 12, we see that active M-dwarfs (red triangles) lie directly along the MS of our training data in both a CMD and all color-color spaces. Thus, we find that only three systems ($3\%$) are classified as high-probabilty WD+MS candidates with our SVM. However, despite these low numbers, the WD+MS candidates that lie close the ZAMS (approximately 25-30% of our final sample) do share some photometric properties with active M-dwarf stars. For these systems, we conclude that some contamination is possible (especially since single M-dwarfs are expected to be more prevalent than WD+MS binaries) and spectroscopy is required to distinguish binarity from potential flaring. In particular, we note (i) Pleiades-c1 is located in this region of parameter space, but is a spectroscopically confirmed binary system and (ii) both Pleiades-c2 and Pleiades-c3 were previously classified as flaring M-dwarfs (Section 6.2.2). However, also note that the presence of M-dwarf flares is not mutually exclusive with a potential WD+MS binary, as post-CE systems with active M-dwarf secondary stars have been identified (e.g. Muirhead et al., 2013).
@@ 6.3.4 Single WDs
Depending on system age, single WDs can have effective temperatures between $10,000K<T_{eff}<40,000K$ (Eisenstein et al., 2006; Gentile Fusillo et al., 2021). Hence, it is possible that older, cooler and redder WDs could share similar photometry to a WD+MS binary. To investigate this potential overlap, we consider the Gentile Fusillo et al. (2021) catalog of WDs identified in Gaia Early Data Release 3. While the full sample contains $\sim 359,000$ WDs, we find that upon cross-matching, only 88 of these systems have 2MASS data.
In Figure 12, we observe that these 88 single WDs fall bluewards of the previously investigated populations in both a Gaia CMD and all color-color spaces. While single WDs appear to occupy extensions of the WD+MS binary population in each color-color space, overlap with WD+MS binaries is present especially when 2MASS photometry is included. Due to this overlap, $38\%$ of the WDs in this catalog are classified by our SVM as probable WD+MS binaries. However, none of our identified WD+MS candidates visually overlap with the single WDs in any parameter space, making it unlikely that any of our candidates are actually single WDs.
Figure 13: The proposed effect of a CE natal velocity kick on the spatial position and kinematics of candidate NGC1342-c1. Left: Spatial location of NGC1342-c1 (star) relative to high-probability cluster stars from Cantat-Gaudin et al. (2020) (gray). Each circle represents the distance NGC1342-c1 would be located if it received a natal kick for various kick magnitudes and travel times. Right: Similar to the left panel, however with proper motion. Each circle here represents the change in NGC1342-c1’s proper motion ($\Delta\mu$) upon receiving a velocity kick completely imparted into the tangential direction (for simplicity).
@@ 6.4 Implications for Natal Kicks
Upon examining the kinematics of the candidate WD+MS systems in our catalog, we find that while their overall proper motions are consistent with their host clusters, the spatial locations for more than half of the candidates are largely offset (see Section 4.2). Thus, we explore whether these spatial offsets could be explained by invoking a natal kick. As a fiducial example, we consider the WD+MS binary candidate NGC1342-c1. NGC1342-c1 has a proper motion consistent with being a member of NGC 1342 ($\chi^{2}=1.03$) but a spatial location inconsistent with the cluster ($\chi^{2}=112.41$) as computed in Section 4.2.
@@
While specific cluster proper motions, distances and escape velocities determine the magnitude to which a natal kick might effect a post-CE system, this analysis shows that a post-CE system can be spatially offset from a cluster despite having a consistent proper motion. Hence, we conclude that it is possible for post-CE systems to migrate far from their birth cluster if a velocity kick is received. Importantly, we emphasize that $78\%$ of the OCs searched in this study revealed no WD+MS candidates. It is almost certain that there exist more WD+MS post-CE cluster binaries than presented in this catalog, however they are likely located outside the search radii chosen in this study. While extending search radii might help identify more WD+MS candidates, finding binaries that have travelled far beyond their host cluster is challenging. Although chemical tagging and dynamical simulations are good tools for associating escaped stars and binaries with individual star clusters (e.g. Grondin et al., 2023, 2024), ample chemistry and radial velocities would be required for definitive associations of these escaped systems.
@@ 7 Summary and Conclusions
In this paper, we search for WD+MS binaries in 299 OCs in the Cantat-Gaudin et al. (2020) catalog. Using photometry from Gaia DR3, 2MASS and Pan-STARRS1, we train a support vector machine model to identify systems that are photometrically similar to a sample of SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010). Throughout, we design both our training sample and method to try to minimize contamination (and thus sacrifice some amount of completeness) in order to identify a set of high-probability candidates, which can then be followed-up for confirmation and characterization. While the focus of this study is identifying WD+MS binaries in clusters, we emphasize that this methodology could be utilized for photometric identification of a variety of other astrophysical sources. Figure 1 highlights the main steps in our identification method. Our main results are summarized below.
@@
    This catalog contains 52 candidate WD+MS binaries in 38 OCs which were identified with high-probability ($P>0.9$) from our SVM classification model and subsequently vetted for data quality (Figure 8). Generally, our catalog contains systems that are redder than the overall sample of WD+MS binaries presented in Rebassa-Mansergas et al. (2010). This is in line with other WD+MS systems identified in Gaia (Inight et al., 2021; Rebassa-Mansergas et al., 2021). 78% of the 299 OCs searched had zero high-probability WD+MS binaries identified by our method.
@@
    We present follow-up spectroscopy from the Lick and Gemini Observatories for three example candidates that span a range of Gaia $BP-RP$ colors (Figure 10). From this spectroscopy alone, it is clear that a variety of systems are present in this catalog. For instance, two objects (which represent both the bluest and reddest systems in our sample) are well-matched to M-type MS stellar models at red wavelengths, but show evidence of broad Balmer lines and excess flux at blue wavelengths—as would be expected for a WD companion. In contrast, the third object shows no clear signatures of a WD in its optical spectrum. It may represent either a case with a limited flux contribution from the WD in the optical regime, or contamination in our sample from an active (flaring) M-dwarf. SEDs for all WD+MS candidate binaries are similar to the three examples shown. Hence, we expect that this catalog spans a wide range of WD and MS configurations.
@@
    Since we solely use photometry to identify WD+MS candidates, we investigate the possibility that other astrophysical sources are present in our catalog. Specifically, we run populations of cataclysmic variables (Inight et al., 2023), MS+MS binaries (Pickles, 1998), rapidly rotating (active) M-dwarfs (Kado-Fong et al., 2016) and single WDs (Gentile Fusillo et al., 2021) through our SVM. Based on this and the locations of our candidate systems in the Gaia CMD and various color-color spaces, we conclude that we are unlikely to have any significant contamination from single WDs or MS+MS binaries. In contrast, our sample may contain CVs in addition to detached WD+MS binaries, and we may have some contamination from young, active M-dwarfs in the $\sim 25-30\%$ of our sample that lies close the the ZAMS. Ultimately, follow-up spectroscopy will be required to fully confirm the nature of each candidate.
@@
Ultimately, this catalog is a necessary first step in a larger effort to provide observational constraints on the CE phase and detailed characterization of a subset of candidates (e.g. Alessi12-c1, Pleiades-c1, etc.) identified in this study is currently underway. Once a larger sample of post-CE binaries in clusters is confirmed, pre-CE progenitor masses can be estimated, leading to a one-to-one mapping between the initial and final masses of systems that have undergone a CE event (similar to how the IFMR for isolated WDs is determined). With these observational benchmarks, this sample will aid in efforts to unlock important new insights into one of the most uncertain phases of binary evolution.
@@ Acknowledgements
The authors thank Max Moe for providing insightful discussion about potential contaminant sources that aided in an improved understanding of the WD+MS candidates in our catalog. The authors also thank Alexander Laroche for useful input in validating our machine learning model and Fraser Evans and Phil Van-Lane for valuable edits in the manuscript. Helpful conversations with Samantha Berek, Victor Chan, Mairead Heiger, Marten van Kerkwijk, Gustavo Medina-Toledo and Jeremy Webb also greatly improved this research.
@@ 8 Data Availability
Data from the Kepler/EPIC (STScI, 2016), TESS Input Catalog and Candidate Target List (STScI, 2018) and Pan-STARRS DR1 Catalog (STScI, 2022) were accessed from the Mikulski Archive for Space Telescopes (MAST) database.
@@ Appendix A Photometric Search Examples
Here, we present results from the SVM classification for two example OCs: Alessi 12 and NGC 2682. As seen in both Figures 14 and 15, the majority of sources in the Alessi 12 and NGC 2682 stellar samples lie along/near the ZAMS in both sample color-color spaces and a Gaia CMD. While the SVM does not identify any probable WD+MS candidate binaries in NGC 2682, one high-probability source ($P>0.9$) and two medium-probability sources ($P>0.5$) are found in Alessi 12.
@@ Appendix B Photometric Errors of High-Probability Candidates
In Figure 16, we present the photometric errors of all 52 high-probability candidate WD+MS binaries in this catalog. Overall, the photometry (including error bars) spans the full WD+MS sample, with only small overlap with the ZAMS. We note that the largest photmetric errors arise from 2MASS $J$ and $K$ bands.
Figure 16: Photometric errors of the high-probability candidate WD+MS binaries in this catalog.
@@ Appendix C Spectral Energy Distributions
As described in Section 3.7, we perform additional vetting of our high-probability candidates through a visual inspection of candidate SEDs. In Figures 17 and 18, we present SEDs for all 52 high-probability candidate WD+MS systems in our catalog. We include both Pan-STARRS $g$, $r$, $i$, $z$ and $y$ band along with 2MASS $J$, $H$ and $K$ band photometry. To guide the eye, we have included five MS Pickles models (Pickles, 1998) spanning K0V to M5V spectral types. However, we emphasize that the observational SEDs have not been corrected for Galactic extinction.
Figure 17: Spectral energy distributions of the high-probability candidate WD+MS binaries in this catalog.